### v2

# CogME: A Cognition-Inspired Multi-Dimensional Evaluation Metric for Story Understanding

Abstract
We introduce CogME, a cognition-inspired, multi-dimensional evaluation metric designed for AI models focusing on story understanding. CogME is a framework grounded in human thinking strategies and story elements that involve story understanding. With a specific breakdown of the questions, this approach provides a nuanced assessment revealing not only AI models’ particular strengths and weaknesses but also the characteristics of the benchmark dataset. Our case study with the DramaQA dataset demonstrates a refined analysis of the model and the benchmark dataset. We argue the need for metrics based on understanding the nature of tasks and designed to align closely with human cognitive processes. This approach provides insights beyond traditional overall scores and paves the way for more sophisticated AI development targeting higher cognitive functions.
Keywords: artificial intelligence; video story understanding; video question answering; evaluation metric
Introduction
In recent years, the development of artificial intelligence (AI) models, particularly pre-trained Large Language Models (LLMs) (?, ?) and diffusion models (?, ?) has made remarkable progress. These models have demonstrated impressive performance in creative tasks, including generating images (?, ?), videos (?, ?), and narrative storytelling (?, ?).
However, skepticism remains about their ability to ‘understand’, as highlighted in recent discussions (?, ?, ?). Indeed, while AI models often outperform humans in generating text and images, their performance in understanding does not reach their outstanding generative outputs (?, ?). This limitation becomes prominent in multi-modal AI models, where integrating and interpreting various data forms – such as image, video, and textual information – imposes considerable challenges. In line with this limitation, video story understanding models demonstrate a significant gap compared to human story comprehension (?, ?).
Another key issue is the inadequacy of current evaluation metrics for AI models. There are widespread arguments that existing metrics are too general and fail to thoroughly analyze the performance of these models. These metrics often rely on aggregate scores, which can obscure the true performance of a model and hinder an understanding of the benchmark dataset’s detailed features used for training (?, ?, ?). This issue becomes increasingly pronounced in more complex tasks, like video story understanding. Consequently, there is an urgent need for a new evaluation method that aligns more closely with human cognition for a more fine-grained assessment of AI model performance. Addressing this pressure would require a thorough understanding of the intrinsic nature of the given tasks and the target-oriented evaluation.
Figure 1: An illustration of CogME framework for an example of DramaQA dataset. It shows a situation in which an Agent predicts the Answer from the given Video clip and Question. Orange arrows indicate the process involves three story understanding components: TARGET, CONTENT, and THINKING.
In light of this need, we have proposed a novel evaluation metric for the story understanding model, named Cognition-inspired Multi-dimensional Evaluation (CogME). CogME is designed to evaluate story understanding AI model performance through a unique lens: A specific breakdown of the questions. The breakdown is grounded in multi-dimensional criteria considering human thinking strategy and story elements.
The unique design is based on the following proposition: If an agent answered a specific question appropriately, it means that “The agent understood the CONTENT of the TARGET through a way of THINKING.” It also means “The question required the agent’s THINKING about the CONTENT of the TARGET.” From the proposition, our approach provides a richer explanation of how well AI models handle different queries by intricately analyzing the context of these questions. This framework not only evaluates the AI model in terms of strengths and weaknesses in responding to various queries but also delineates the specific features of the dataset.
We also present the results of an in-depth case study using DramaQA, a representative benchmark dataset for video story understanding, along with its baseline model. Fig. 1 illustrates how we apply the CogME to an instance of DramaQA. Our results demonstrate that using CogME allows for a more thorough and systematic evaluation of both the benchmark datasets and the AI models.
Related Work
Narrative Comprehension of Human
Cognitive science studies on narrative comprehension
, conducted through various reading comprehension tasks, have yielded seminal findings. It was found that people tend to prioritize the main aspects of a story over individual parts, indicating a higher level of cognitive engagement with narrative comprehension (?, ?). The concept of schema (?, ?, ?) in memory reveals that recall is an active process influenced by personal and cultural contexts, emphasizing the critical roles of context and prior knowledge in comprehension (?, ?, ?). Mental representations such as scripts and schemata were denoted as crucial cognitive structures aiding comprehension and inference-making (?, ?, ?).
From these research bases, the Situation model and Event-indexing model enlightened the analytical approach to human story understanding and established the foundation of structured evaluation. Situation model constructed detailed mental models during story comprehension as encompassing events, characters, and settings (?, ?). The event-indexing model puts the emphasis on human understanding that is divided into five independent dimensions when reading: time, space, character, causality, and motivation in the context of the narrative (?, ?).
Research on understanding video stories
is relatively scarce compared to enriched studies on reading comprehension despite the rapidly increasing importance of video narratives in our daily lives. Since reading and viewing are distinct tasks with different cognitive loads (?, ?, ?), it is undesirable to replicate reading comprehension structures for video comprehension (?, ?, ?).
Recently, there have been research efforts focused on understanding video narratives within real-world contexts. Specifically, these studies employ a top-down approach to observing brain activity not in controlled or fragmented experimental videos, but rather in watching typical movie scenes (?, ?, ?). These studies contrast with conventional cognitive psychology studies, which mainly rely on a bottom-up approach focused on the segmented form of visual and audio stimulations (?, ?). Therefore, it is necessary to take an integrated approach to fully grasp how individuals engage with and interpret video narratives in real-world situations, encompassing both the perceptual and narrative aspects of video understanding.
Evaluation of Machine Comprehension
Machine Reading Comprehension (MRC)
is the most prominent area regarding machine comprehension (?, ?). Question-answering (QA) has been widely adopted to evaluate text understanding of AIs (?, ?, ?). However, efforts to reflect the content structure of the text to assess the MRC model have been made only recently. Dunietz et al. argued that the insufficient clarity of existing metrics for MRCs could be improved by applying appropriate templates derived from the definition of comprehension (?, ?). To address this issue, they employed the Event-indexing model from human studies (?, ?) to posit four elements that machines should incorporate for better reading comprehension: place, time, causality, and motivation. Similarly, Weston et al. demonstrated empirical test results of multiple AI agents’ textual understanding abilities with structured QA skillsets categorized twenty types of questions for understanding and reasoning with text (?, ?). However, these tests are limited to being constructed with fragmentary and artificial descriptions. Further research is needed before their method can be applied at the level of complexity and richness typically found in everyday human storytelling.
Video understanding models
have yet to keep pace with the growing demand, even though video-based storytelling has recently emerged as one of the most prominent forms of media content. Developing AI that can understand video stories is challenging, given that it requires an all-inclusive process to analyze images, scripts, and sounds with temporal dependencies, natural language, and various levels of reasoning (?, ?). Despite the difficulty, several efforts to develop video understanding AI have centered on large-scale video datasets (?, ?, ?, ?, ?, ?, ?), but the technologies do not extend beyond simple image processing tasks such as detecting or tracking objects.
From the evaluation perspective, the most prevalent approach involves building massive QA datasets, leveraging open-ended or multiple-choice QAs for AI training and testing (?, ?). However, existing evaluation methods heavily count on unidimensional metrics (?, ?), such as basic QA accuracy scores for the multiple-choice QA datasets, which often fall short in providing a comprehensive explanation of the model’s performance. In the case of open-ended QA, the automatic evaluation primarily depends on n-gram-based sentence similarity measures such as BLEU (?, ?), METEOR (?, ?), ROUGE (?, ?), and CIDEr (?, ?). These metrics, focusing on the similarity between the reference and generated sentences, frequently do not align with human judgments, reflecting a significant disconnect. These automatic metrics are usually insufficient to provide a refined interpretation like the strengths and weaknesses of the AI model (?, ?).
On the other hand, there are several efforts in human evaluations to appraise the effectiveness of automatic metrics in accurately assessing AI performance. (?, ?, ?). While these efforts indicate that human evaluations are beneficial for assessing AI performance, they also reveal a significant challenge: the fine-grained evaluation is not effectively achieved in proportion to the resources and costs involved in the evaluation process.
Accordingly, evaluating AI solely based on its QA accuracy or similarity to human performance is inadequate, especially for tasks with high complexity, such as video story understanding. In evaluating the performance of video understanding AI, there arises a need for a structured framework that meticulously analyzes both the nature of the understanding process and the unique characteristics of the medium. In response to this, we propose a new metric that not only reflects the existing frameworks of human story comprehension but also incorporates the distinct features of video storytelling.
Table 1: The sub-components within TARGET
Elements Definition
Character Information of individuals featured in the video.
Object Items and body parts featured in the video.
Place Spatial information of the story in the video.
Conversation Characters’ dialogues, monologues, speech sounds, and text messages.
Behavior Movements and actions of the subjects in the video.
Event Information about what happened in the video.
Emotion The feelings expressed by the subject in the video.
Commonsense Concepts and knowledge that people universally accept in a given culture.
Table 2: The sub-components within CONTENT
Elements Definition
Identity Personal information of subjects or names of objects in the story.
Feature Characteristics, traits, or atmosphere of subjects and/or objects.
Relationship The relationships between two or more targets.
Means Instruments or methods used to achieve a particular purpose.
Context Storyline revealed through the conversations or interactions between characters.
Sequence Related events with time series and the changes before and after.
Causality Causes and consequences of a particular change: natural or mechanical.
Motivation Changes resulting from actions involving personal preferences or intentions.
Table 3: The sub-components within THINKING
Elements Definition
Recall Retrieving or recollecting factual information in the scene.
Grasping Perceptions or interpretations of the scene with temporal and spatial changes.
Reasoning Making logical judgments from circumstantial evidence not direct observations.
New Evaluation Paradigm Based on the Understanding Processes of Humans
To evaluate understanding competence thoroughly, we developed a multifaceted paradigm integrating video narrative elements and thinking strategies involving the queries. In analyzing the story elements provided by the video, we have adapted the Situation Model (?, ?) and the Event-Indexing Model (?, ?) and expanded them to better suit video narratives. Regarding human thinking strategies, we referred to Bloom’s Taxonomy, widely accepted as a representative framework demonstrating the hierarchical structure of cognitive processes (?, ?, ?). Each level of Bloom’s taxonomy represents a different cognitive skill, ranging from the basic recall of facts and grasping details to reasoning hidden matter and, ultimately, evaluation and creation.
Figure 2: Examples of tags applied to questions a Cases of tagged to questions in shot-level video, which require simple recall. b Cases of tagged to questions in scene-level video, which require comprehensive reasoning.
Drawing on insights from previous models in cognitive science, we have developed a multi-dimensional metric, CogME, which consists of three key components: TARGET, CONTENT, and THINKING. TARGET refers to the information perceived by watching the video, CONTENT to the knowledge acquired through the target information, and THINKING to the cognitive process of deriving knowledge from the information. The understanding procedure involves integrating the three components, representing cognitive processes, to interpret the story elements presented in the video. This integration can be succinctly expressed as “I understand the CONTENT of the TARGET through a way of THINKING.”
The sub-components of TARGET and CONTENT were determined by analyzing the story elements necessary for understanding video stories. The sub-components of THINKING were decided to align with the range required among human thinking strategies. Tables 1 – 3 demonstrate the sub-components of TARGET, CONTENT, and THINKING derived from story elements and thinking strategies.
Materials and Methods
Application to VideoQA Dataset: DramaQA
[…]
Annotating the Understanding Components
To determine what information, knowledge, and thinking strategies are required for answering questions, we annotated 4,385 questions from episodes 13th to 15th using the understanding sub-components defined in the CogME framework. Two specialists in cognitive science conducted an elaborate analysis of the given videos and the questions and tagged the required sub-components in each question.
Fig. 2 illustrated two examples of this annotation. [^1] The DramaQA dataset features memory capacity criteria related to the length of the video segments. The criteria include two types of video clip: 1) shot-level video without camera cut, spanning a few seconds, and 2) scene-level video with multiple events in a single location, spanning a few minutes Fig. 2(a) shows an example of a question answerable by simply recalling a single cue from a shot-level video. It involves identifying clothing-related information about the only person in the shot. In contrast, Fig. 2(b) illustrates a more complex scenario where annotating sub-components in a scene-level video highlights a different level of narrative understanding from the example shown in Fig. 2(a). During a three-minute and six-second runtime, two characters are shown walking in the park and chatting. To answer the question, the agent must not only recall and grasp the content of the conversation but also infer why Heeran said to envy Haeyoung, which is not directly mentioned in the dialog.
For the THINKING strategies, we assumed that the higher cognition embraces the lower ones based on the hierarchical Bloom’s taxonomy (?, ?). Therefore, we only labeled the highest thinking component. However, regarding the TARGET and CONTENT of the question, we tagged up to three if multiple sub-elements were involved in a single question.
Figure 3: Performance profiles of the two models. The vertex of each polygon represents the ratio (%) of correct predictions for the DramaQA dataset. Each radar plot represents TARGET(left), CONTENT(middle), THINKING(right) component. Light blue areas indicate the performance profiles of Agent I (MCM model), and pink areas display the performance profiles of Agent II (MemN2N baseline model).
Scoring the Questions and Prediction Results
In the context of the logical complexity of the THINKING module, a weight of 2 and 3 was assigned to grasping and reasoning each, assuming grasping includes the simple recall and the reasoning process retains the recall and grasping.
Even identical questions can be endowed with different weights. This variance depends on the diverse contexts of the given video, ultimately affecting the complexity of the question. Considering the question, ‘Who is smiling?’, if there is only one character in the video, the prediction can be made simply by recalling who it is. However, when there are two or more characters, that prediction requires a higher level of thinking. This involves discriminating the smiling character from the others and then identifying that character. This grasping process certainly includes recalling relevant details.
All tagged sub-components were multiplied by the weight given to the thinking component, assuming that the degree of difficulty of the thinking strategy determines the difficulty of the entire question. In the example shown in Fig. 2, for question in Fig. 2(a), which requires recall, all labeled elements are worth 1 point, while for question in Fig. 2(b), which requires reasoning, all labeled elements are worth 3 points.
The model’s accuracy for each sub-component was calculated by scoring correct predictions as 1 and incorrect ones as 0. We then computed the overall success rate by comparing the number of correct predictions to the total attempts. To address the imbalanced frequency of each sub-component, the achievement rate was expressed as a percentage.
The analysis was conducted in R 4.1.1 (?, ?) environment for arithmetic computing and visualization. The polygonal profiles were produced using the fmsb 0.7.1 package (?, ?).
Comparison of Multiple Agent Performances
In additional experiments, two agents trained with the same dataset were evaluated with CogME to compare their performances. The two models were the MCM model (?, ?) (Agent I) and the baseline model of MemN2N (?, ?) (Agent II) [^2] It should be noted that these experiments were not meant to identify the model with the better performance but only to compare them objectively. Accordingly, in the results section, we refer to the models as Agent I and Agent II instead of their names. The two models were examined after being trained on the same dataset, DramaQA. Their performance profiles were generated using CogME.
Results
Evaluating Model Performances
Fig. 3 shows the accuracy profile for each understanding component obtained from applying CogME to the dataset. The profile indicates that the Agents demonstrate varying levels of competence across different sub-components, as depicted by the uneven shape of the polygon.
Additionally, we observed distinct disparities when comparing the results of two different models, Agent I (represented by a light blue polygon) and Agent II (represented by a pink polygon), both trained on the same dataset. The overall correct prediction rates were 73.4% for Agent I and 58.7% for Agent II, indicating a difference of 14.7%. However, as detailed in Fig. 3, this discrepancy is not uniform across all areas but varies by specific factors. For instance, Agent II leads by approximately 1% in questions requiring the identification of the Means, while the gap extends to 20% in questions involving the Conversation. Such variations highlight the importance of a multi-dimensional evaluation approach to assess the capabilities of AI models accurately.
Analyzing Questions in DramaQA Dataset
Figure 4: Frequencies of sub-components tagged in the questions of the DramaQA dataset. Each bar shows the number of times a sub-component was labeled out of 4,385 questions.
Alongside the performance profiling, annotations based on CogME enable us to figure out the dataset’s features in terms of data distribution. Fig. 4 shows the distribution of sub-components tagged for the questions in the DramaQA dataset, which characterizes the dataset. The analysis reveals that the distribution of sub-components is uneven. Moreover, the sub-components significantly underrepresented in the dataset tend to align with lower accuracy in model performance. Notably, all four elements that appeared less than 5% frequently in the dataset (i.e., Commonsense, Relationship, Means and Causality) showed low accuracies that are below 50%; Commonsense (35.8%), Relationship (47.8%), Means (34.3%), Causality (47.2%).
Discussion
In this study, we introduced a novel framework, CogME, which is centered around the features of the posed questions. This approach is grounded in structured metrics considering human thinking strategy and story elements using a top-down approach. Unlike conventional AI evaluation methods, which emphasize overall scores that lead to a lack of robust evaluation, CogME provides a multi-dimensional quantified profile for AI models. This profile provides insight into the model’s strengths and weaknesses in understanding abilities by applying the metric to an existing dataset.
When we compared the two models according to the CogME metric, their detailed dissimilarities were revealed, which could not be explained by the aggregated scores. It provided thorough assessments and comparisons of each AI model’s understanding abilities. This assessment is expected to be applicable regardless of whether the agent is a machine or a human and also provide a comprehensive quantification of the features of the agents’ levels of understanding.
Furthermore, CogME’s fine-grained evaluation approach not only assesses the AI models thoroughly but also offers a comprehensive analysis of the benchmark dataset, thus providing a more in-depth understanding of the models’ capabilities. An instance of this analysis, the relationship between low frequency and low accuracy [^3] Although we only analyzed questions from the validation set, we assumed that the CogME profile of the training set would be similar based on the preliminary analysis that demonstrated a similar distribution of question types across the datasets. (?, ?), suggests that a learning deficiency can be a deciding factor in aggravated QA performance. As mentioned in the ‘Narrative Comprehension of Human’ section, humans tend to focus on the central aspect of the story rather than on individual instances (?, ?). However, our analysis reveals that the breakdowns of this dataset predominantly collect fragmentary information, rather than focusing on the central or structural aspects of the story. This observation suggests that a dataset design that overlooks the task’s fundamental aspects could hinder the development of AI aimed at higher cognitive functions.
It would also be worth establishing a CogME framework suitable for setting data characteristics in advance when designing new datasets. As Fig. 4 shows, when generating a massive QA dataset, it was difficult to regulate sufficient variety in the types of questions or the sub-components. These maldistribution issues have been noted not only in DramaQA but also in many other QA datasets (?, ?). In this context, this primary setup could work as a theoretical foundation to ensure proper data allocation in the dataset, whether through crowdsourcing or automatic question generation.
A hurdle of our study is that the sub-components encompassed in the provided questions and videos were manually annotated. This was an inevitable choice to comprehensively reflect the elements because even the same questions can have different contexts in the given video (see ‘Scoring the Questions and Prediction Results’ section). Although manual annotation remains quite an obstacle in utilizing CogME, it ensures that the evaluation accurately encompasses the diverse aspects of the dataset, aligning closely with the intricate nuances of both the video content and the associated queries. However, in the future, we may consider using a multi-modal classification model that automatically annotates the sub-components in CogME for given questions and related video sequences. Such automation would not only simplify the evaluation process but also allow the scalability of applying these methods to larger and more complex datasets.
Moreover, scoring for multiple-choice questions can result in the loss of some information. According to our annotating, even if the model correctly recognized the Identity of Character, if it failed to infer other information, such as Emotion, the Character’s score for this question would be zero. However, by integrating it with methods akin to rubrics emphasized in pedagogy (?, ?), we argue that this metric has the potential to be extended to other tasks such as open-ended or fill-in-the-blank tests, summaries, and rewriting, which can be analyzed using understanding sub-components.
In conclusion, this study introduces the CogME framework, which provides a multi-dimensional analysis focusing on story understanding. Our approach reveals deeper insights, surpassing the limitations of overall scores like correct answer rates. The potential of CogME extends to various AI tasks and dataset designs, suggesting its adaptability and utility in advancing AI assessment toward more nuanced and sophisticated dimensions. This study sets a new benchmark, paving the way for more comprehensive approaches to developing AI agents.
Acknowledgments
Anonymous CogSci submission. This will be updated in the final version.